# An Empirical Study on Compliance with Ranking Transparency in the Software Documentation of EU Online Platforms

### 22 Dec 2023 v1

@@ 6.2. Large-scale Manual Assessment
Our larger scale empirical study involved 134 non-expert participants, sourced through Prolific. [^7] https://www.prolific.co Following the same methodology as the study conducted with experts (Section 6.1), participants were asked to evaluate the documentation identified for Tripadvisor and Booking presented in Section 4.1. They rated, on a 1-5 scale, how effectively the documentation answered questions related to the explanatory information mandated by the P2B regulations. Guided by the European Commission guidelines (Section 3.1) stating that explanations about the ranking system should cater to the ‘average’ users’ technical aptitude and needs for a given service (Section 3.1), participants were instructed to adopt the perspective of business users employing online intermediation services or online search engines. To maintain the study’s feasibility and ensure a median completion within approximately 10 minutes, we considered only four of the 17 questions presented in Section 4 and limited the analysis to the documentation of two intermediation services: Booking and Tripadvisor. These were chosen based on the brevity of their explanations (Booking is seven webpages long, Tripadvisor nine) and were representative of the best and worst explanations among the selected set of online intermediation services, as adjudged by the legal experts. We curated the four questions, drawing from the list in Section 4, by considering the objective requirements of the regulation and its main goal, i.e., being able to understand how the ranking works and understanding how to improve/change the results of the ranking to improve outcomes. Eventually, the chosen questions rephrased so as to be compatible with a 1 (strongly disagree) to 5 (strongly agree) scale evaluation, were:
@@ 4. Q16
Figure 1. Average agreement rates for participants reviewing documentation on Tripadvisor and Booking based on different time durations. It showcases results considering minimum review durations of five to zero minutes. We also show statistically significant p-values comparing agreement rates between the two platforms for the same duration.
The study was designed to have evenly distributed male and female participants. Participants were only allowed from the UK [^8] Since after Brexit, the UK has the discretion to retain, modify, or discard the provisions of the P2B Regulation. or Ireland which are the only English-speaking countries where the P2B Regulation applies. We applied further pre-screening on the spoken language (participants had to be fluent in English) and the Prolific approval rate (Eyal et al., 2021) which had to be above 75%. Eventually, the study results aligned with our initial expectations, with observed differences between the perceived quality of explanations in the documentation of Tripadvisor and Booking. Furthermore, a few one-sided Mann-Whitney U tests also confirmed that these differences are statistically significant in most cases, as summarized in Figure 1:
@@
As shown in Figure 1, a deeper engagement with the documentation—measured by spending more minutes—highlighted the differences between Booking and Tripadvisor. Intuitively, the longer one spends reading the documents, the more likely they are to notice the nuances and distinctions. However, only 55 of the 134 participants dedicated more than 5 minutes to reading the documentation. Moreover, the data in Figure 1 closely align with the experts’ evaluations for each question. For instance, for Q3, experts ranked Tripadvisor at an average of 4.66, a score significantly higher than Booking’s 2.33. In Q4, experts assigned Tripadvisor a rating of 3.66, outperforming Booking’s 1.66. During Q15’s assessment, Booking garnered a score of 4, marginally edging out Tripadvisor’s 3.33. For Q16, experts gave Tripadvisor a 2.66, with Booking trailing at 1. By cross-referencing the figures and experts’ scores, there is a consistent pattern between the two, reinforcing the validity and reliability of the study findings. Importantly, these findings were not only statistically significant but their effect sizes were medium to large, according to Cohen’s conventional criteria. In particular, the effect sizes for Q3 and Q4 demonstrate a moderate practical advantage for Tripadvisor. While a large effect size for Q15 underlines the shortcomings of Tripadvisor when compared to Booking.
@@ 6.3. Automated Assessments
As demonstrated so far, online platforms often struggle to produce clear documentation that complies with regulations like P2B. Evaluating and ensuring the quality of such documentation is challenging, largely due to the lack of standardized automated tools for assessing compliance and explainability across different platform providers. To address this, we employed the assessment tools detailed in Section 5. Our aim was to determine if these tools’ evaluations are consistent with expert opinions and the empirical findings discussed in Sections 6.1 and 6.2. We also compared their performance against baseline methods such as constant ”Yes” responses and random [^9] We used 42 as a random seed. assessments. The findings presented in Table 4 reveal that both the DoX-based method and ChatGPT 3.5 align with the experts’ evaluations 63.88% of the time. However, ChatGPT 4 ranks as the least effective AI-based tool, under performing even the random assessment method. The scoring between the experts and ChatGPT was normalized as follows: any response scoring 3 or higher (indicating the question is “partially answered”, as detailed in Section 6.1) was categorized as ‘Yes’, while scores below 3 were labeled ‘No’. The DoX-based method already provides a straightforward ‘Yes’ or ‘No’ response.
@@
We also performed Mann-Whitney U tests and a rank biserial correlation analysis to analyze the alignment between the explanatory relevance, pertinence, and DoX scores produced by the second tool (Section 5.2) and the yes-or-no experts’ answers. As shown in Figure 2, all three scores show a significant $p$-value (less than .05), implying that the difference between the ‘Yes’ and ‘No’ groups in experts’ answers is statistically significant for each score. The rank biserial correlation values range from -.456 to -.539, thus suggesting a moderate to strong negative correlation between the scores and the experts’ answers. The explanatory relevance score correlates the best with the majority of answers, indicated by the highest absolute value of the rank biserial correlation (-.539).
@@ 7. Discussion & Threats to Validity
The findings from Sections 6.1 and 6.2 highlight a statistically significant disparity in compliance scores across diverse platforms. This implies varying levels of compliance among platforms regarding ranking transparency. A possible reason for this inconsistency is the absence of standardized automated tools for gauging compliance and explainability. Section 6.3 implies that automated tools could offer a solution. Notably, results indicate that ChatGPT 3.5 and the DoX-based approach, the best ones, align with the majority of experts 63.88% of the time. However, they are far from flawless, indicating there is considerable potential for refinement in automated assessment technologies. Delving deeper, the DoX-based approach seems to perform better than GPT-3.5 when evaluating longer documents. As illustrated in Table 1, for documentation from Google (52 docs, averaging 1679 words each), Bing (16 docs, averaging 964 words each), and Tripadvisor (10 pages, averaging 1653 words each), the DoX-based method achieves agreement rates of 52.63%, 68.42%, and 64.71%, respectively. In contrast, GPT-3.5 scores 36.84%, 68.42%, and 58.82%. These results can be attributed to the design of the DoX-based approach, which is optimized to handle long documents effectively, bypassing the input size constraints of ChatGPT. Our study used ChatGPT without any fine-tuning. While the DoX-based approach incorporates GPT-4, its performance surpasses that of GPT-4 without answer retrieval. This hints at the instrumental role that prompt engineering might have in bolstering these tools’ efficacy. It is also crucial to understand the nuances of the DoX-based approach. Though this approach aligns with experts 63.88% of the time, this alignment merely considers ChatGPT’s binary (‘yes’ or ‘no’) responses and omits the explanatory relevance scores. An in-depth examination of the alignment between the explanatory relevance scores from our DoX-based methodology and expert responses reveals their potential in pinpointing documentation sections that are non-compliant or inadequately articulated. Figure 2 illustrates that lower scores correlate with expert disapproval. Thus, the feedback depth from the DoX-based approach exceeds that of ChatGPT. We acknowledge a potential limitation of this study: As discussed in Section 4.1, the study focused solely on textual content. However, some platforms also feature video content or imagery that could provide additional insights. Yet, for intermediation services at least, it is compulsory to provide their explanations in written form (Section 4.1), hence our decision to focus on textual content. We cannot rule out that information not available in textual content could be embedded in videos or images. Given the performance scores observed, it is clear that the fully automated tools we have introduced cannot and should not entirely replace legal experts. While these tools may be effective at continuously tracking changes in documentation, their best use is in tandem with legal professionals. The data in Table 4 highlights cases in which ChatGPT 3.5 is the top performer, and in some cases, the DoX-based method is more effective. Thus, for the best results, it seems reasonable to suggest combining both tools, then proceeding with manual validation.
@@ 8. Societal Implications
The qualitative assessment approaches to compliance measurement, and automated tools that we proposed in the previous sections have various societal implications, which we underline hereafter. Our study provides concrete data and assessments of major market actors’ documentation, in terms of compliance with a specific piece of EU law. It goes beyond existing literature by combining insights from legal experts, laypersons, and automated tools. This research can highlight ranking practices impacting individuals, businesses, and society. It can also assist platforms in understanding their level of compliance compared to competitors and addressing any issues. These results are beneficial for society. Additionally, the paper highlights that assessing the quality of software documentation is a difficult, resource-consuming task, due to the lack of standardized automated tools to evaluate compliance. The paper also proposes a solution to this gap, in the form of automated assessment tools that are tested against human judgment. The societal benefits of this solution include enhanced means for regulators and public authorities to verify and force compliance with ranking transparency requirements. This, in turn, may lead to more balanced and fair markets for business and corporate website users, as well as more competition between platforms, as intended by the EU legislator. Such societal benefits can help achieve the United Nations Sustainable Development Goal 10.3 (of Economic and Affairs, 2023), which seeks to reduce the inequalities of outcomes and ensure equal opportunities. Finally, our research explores approaches that can measure the quality of explanations provided about software and algorithmic processes vis-à-vis a legal standard, which is actually a crucial field of research. This is particularly due to the rapid development of (i) machine learning techniques and tools, (ii) methods to explain machine learning systems and their outputs (Speith, 2022), and (iii) legal requirements imposing explanations when such systems are used (Hacker and Passoth, 2020). In the context of our paper, it is clear that the ranking systems at stake include machine learning technology, which can be difficult to explain to laypersons. As lawmakers impose on online platforms to disclose information about these systems, it is particularly beneficial to society to be able to measure the quality of the explanations provided and determine when an explanation is insufficient or of poor quality. Besides, this reasoning and the tools that we develop can be transposed to other contexts implying the compulsory provision of explanations in relation to machine learning tools or outputs, such as the forthcoming AI Act (see, for instance, Art. 13 of the proposed AI Act (European Commission and Technology, 2021)).
9. Conclusions and Future Directions
The EU’s P2B Regulation requires online platforms, like intermediation services and search engines, to clearly explain their ranking procedures. In this context, we studied how platforms like Amazon, Booking, Google, Bing, Tripadvisor, and Yahoo comply with this regulation. With three legal experts and feedback from 130 users, we analyzed the documentation. Our findings showed varying compliance levels. One reason might be the lack of standardized assessment criteria or automated tools. To address this, we introduced two automated assessment tools. First, given its popularity, we considered ChatGPT as an assessment method. Next, we proposed a hybrid method (designed to better handle long documentation) merging ChatGPT with answer retrieval techniques and a new metric, DoX, to measure documentation explainability. When we compared the results from our automated tools to manual assessments, they agreed with human evaluations 63.88% of the time. This suggests these tools could help make the review process more efficient. Importantly, the method based on DoX was particularly good at assessing lengthy documents. Furthermore, its explanatory scores aligned well with expert opinions, implying that content with weak explanations tends to receive lower compliance ratings from professionals. While these tools have limitations, their continuous assessment capability can be invaluable for regulators and the public. Combining them with expert judgment can promote transparent documentation, essential for a fair digital business space. This transparency aligns with goals like the UN’s Sustainable Development Goal 10.3, fostering an informed and equitable online ecosystem. Future improvements could refine these tools through better prompt engineering, tailoring models like ChatGPT for compliance tasks, and including multimedia analysis capabilities.
Declarations
@@ 9. Conclusions and Future Directions
Acknowledgements. F. Sovrano and A. Bacchelli gratefully acknowledge the support of the Swiss National Science Foundation through the SNF Project 200021_197227.
@@
Author Contributions. F. Sovrano & M. Lognoul: conceptualization, methodology, data curation, original draft preparation, investigation, validation, formal analysis. F. Sovrano: software, visualization. A. Bacchelli: review, and editing.

### 23 Jan 2024 v2

@@ 6.2. Large-scale Manual Assessment
Our larger scale empirical study involved 134 non-expert participants, sourced through Prolific. [^6] https://www.prolific.co Following the same methodology as the study conducted with experts (Section 6.1), participants were asked to evaluate the documentation identified for Tripadvisor and Booking presented in Section 4.1. They rated, on a 1-5 scale, how effectively the documentation answered questions related to the explanatory information mandated by the P2B regulations.
Guided by the European Commission guidelines (Section 3.1) stating that explanations about the ranking system should cater to the ‘average’ users’ technical aptitude and needs for a given service (Section 3.1), participants were instructed to adopt the perspective of business users employing online intermediation services or online search engines.
To maintain the study’s feasibility and ensure a median completion within approximately 10 minutes, we considered only four of the 17 questions presented in Section 4 and limited the analysis to the documentation of two intermediation services: Booking and Tripadvisor. These were chosen based on the brevity of their explanations (Booking is seven webpages long, Tripadvisor nine) and were representative of the best and worst explanations among the selected set of online intermediation services, as adjudged by the legal experts.
We curated the four questions, drawing from the list in Section 4, by considering the objective requirements of the regulation and its main goal, i.e., being able to understand how the ranking works and understanding how to improve/change the results of the ranking to improve outcomes.
Eventually, the selected questions, modified to fit a 1 (strongly disagree) to 5 (strongly agree) scale, were:
@@ 4. Q16
Figure 1. Average agreement rates for participants reviewing documentation on Tripadvisor and Booking, considering review durations from five to zero minutes. Includes statistically significant p-values for agreement rate comparisons between platforms.
The study was designed to have evenly distributed male and female participants. Participants were only allowed from the UK [^7] Since after Brexit, the UK has the discretion to retain, modify, or discard the provisions of the P2B Regulation. or Ireland which are the only English-speaking countries where the P2B Regulation applies. We applied further pre-screening on the spoken language (participants had to be fluent in English) and the Prolific approval rate (Eyal et al., 2021) which had to be above 75%.
Eventually, the study results aligned with our initial expectations, with observed differences between the perceived quality of explanations in the documentation of Tripadvisor and Booking. Furthermore, a few one-sided Mann-Whitney U tests also confirmed that these differences are statistically significant in most cases, as summarized in Figure 1:
@@
As shown in Figure 1, a deeper engagement with the documentation (achieved by spending more minutes) highlights the differences between Booking and Tripadvisor. That is because, intuitively, the longer one spends reading the documents, the more likely they are to notice nuances and differences.
However, only 55 of the 134 participants dedicated more than 5 minutes to reading the documentation. This points to the importance of providing brief and clear summaries at the start of documentation for quick information access. This would ensure that the most relevant information is efficiently communicated to non-expert users, within a reasonably limited timeframe.
The data in Figure 1 closely align with the experts’ evaluations for each question. For instance, for Q3, experts ranked Tripadvisor at an average of 4.66, a score significantly higher than Booking’s 2.33. In Q4, experts assigned Tripadvisor a rating of 3.66, outperforming Booking’s 1.66. During Q15’s assessment, Booking garnered a score of 4, marginally edging out Tripadvisor’s 3.33. For Q16, experts gave Tripadvisor a 2.66, with Booking trailing at 1.
By cross-referencing the figures and experts’ scores, there is a consistent pattern between the two, reinforcing the validity and reliability of the study findings.
Importantly, these findings were not only statistically significant but their effect sizes were medium to large, according to Cohen’s conventional criteria. In particular, the effect sizes for Q3 and Q4 demonstrate a moderate practical advantage for Tripadvisor. While a large effect size for Q15 underlines the shortcomings of Tripadvisor when compared to Booking.
@@ 6.3. Automated Assessments
As shown so far, online platforms often struggle to produce clear documentation that complies with regulations like P2B. Evaluating and ensuring the quality of such documentation is challenging, largely due to the lack of standardized automated tools for assessing compliance and explainability across different platform providers.
To address this, we employed the assessment tools detailed in Section 5. Our aim was to determine if these tools’ evaluations are consistent with expert opinions and the empirical findings discussed in Sections 6.1 and 6.2. We also compared their performance against baseline methods such as constant ”Yes” responses and random [^8] We used 42 as a random seed. assessments.
The findings in Table 4 show that the DoX-based method and ChatGPT 3.5 match experts’ evaluations 63.88% of the time. However, ChatGPT 4 is the least effective AI tool, performing even worse than random assessment. Since ChatGPT is a black-box, understanding why this occurs is impossible to us.
The scoring between the experts and ChatGPT was normalized as follows: any response scoring 3 or higher (indicating the question is “partially answered”, as detailed in Section 6.1) was categorized as ‘Yes’, while scores below 3 were labeled ‘No’. The DoX-based method already provides a straightforward ‘Yes’ or ‘No’ response.
@@
We also performed Mann-Whitney U tests and a rank biserial correlation analysis to analyze the alignment between the explanatory relevance, pertinence, and DoX scores produced by the second tool (Section 5.2) and the yes-or-no experts’ answers.
As shown in Figure 2, all three scores show a significant $p$-value (less than .05), implying that the difference between the ‘Yes’ and ‘No’ groups in experts’ answers is statistically significant for each score. The rank biserial correlation values range from -.456 to -.539, thus suggesting a moderate to strong negative correlation between the scores and the experts’ answers. The explanatory relevance score correlates the best with the majority of answers, indicated by the highest absolute value of the rank biserial correlation (-.539).
@@ 7. Discussion & Threats to Validity
The findings from Sections 6.1 and 6.2 highlight a statistically significant disparity in compliance scores across diverse platforms. This implies varying levels of compliance among platforms regarding ranking transparency. A possible reason for this inconsistency is the absence of standardized automated tools for gauging compliance and explainability.
Section 6.3 implies that automated tools could offer a solution. Notably, results indicate that ChatGPT 3.5 and the DoX-based approach, the best ones, align with the majority of experts 63.88% of the time. However, they are far from flawless, indicating there is considerable potential for improvement.
Delving deeper, the DoX-based approach seems to perform better than GPT-3.5 when evaluating longer documents. As illustrated in Table 1, for documentation from Google (52 docs, averaging 1679 words each), Bing (16 docs, averaging 964 words each), and Tripadvisor (10 pages, averaging 1653 words each), the DoX-based method achieves agreement rates of 52.63%, 68.42%, and 64.71%, respectively. In contrast, GPT-3.5 scores 36.84%, 68.42%, and 58.82%. These results can be attributed to the design of the DoX-based approach, which is optimized to handle long documents effectively, bypassing the input size constraints of ChatGPT.
We used ChatGPT without any fine-tuning. While the DoX-based approach incorporates GPT-4, its performance exceeds that of GPT-4 without answer retrieval. This hints at the instrumental role that prompt engineering might have in bolstering these tools’ efficacy.
It is also crucial to understand the nuances of the DoX-based approach. Though this approach aligns with experts 63.88% of the time, this alignment merely considers ChatGPT’s binary (‘yes’ or ‘no’) responses and omits the explanatory relevance scores. An in-depth examination of the alignment between the explanatory relevance scores from our DoX-based methodology and expert responses reveals their potential in pinpointing documentation sections that are non-compliant or inadequately articulated. Figure 2 illustrates that lower scores correlate with expert disapproval. Thus, the feedback depth from the DoX-based approach exceeds that of ChatGPT.
We acknowledge a potential limitation of this study: as discussed in Section 4.1, the study focused solely on textual content. However, some platforms also feature video content or imagery that could provide additional insights. Yet, for intermediation services at least, it is compulsory to provide their explanations in written form (Section 4.1), hence our decision to focus on textual content. We cannot rule out that information not available in textual content could be embedded in videos or images.
Given the performance scores, the fully automated tools should not replace legal experts. These tools are effective at tracking flaws in documentation, but their best use is with legal professionals. Automated tools could be used to determine if existing documentation shows insufficient compliance, and if so, whether a legal expert is needed to enhance the problematic documentation.
Finally, Table 4 highlights cases in which ChatGPT 3.5 is the top performer, and cases in which the DoX-based method is more effective. Thus, for the best results, it seems reasonable to suggest combining both tools, then proceeding with manual validation.
@@ 8. Societal Implications
The qualitative assessment approaches to compliance measurement, and automated tools that we proposed in the previous sections have various societal implications, which we underline hereafter.
Our study provides concrete data and assessments of major market actors’ documentation, in terms of compliance with a specific piece of EU law. It goes beyond existing literature by combining insights from legal experts, laypersons, and automated tools. This research can highlight ranking practices impacting individuals, businesses, and society. It can also assist platforms in understanding their level of compliance compared to competitors and addressing any issues. These results are beneficial for society.
Additionally, the paper highlights that assessing the quality of software documentation is a difficult, resource-consuming task, due to the lack of standardized automated tools to evaluate compliance. The paper also proposes a solution to this gap, in the form of automated assessment tools that are tested against human judgment. The societal benefits of this solution include enhanced means for regulators and public authorities to verify and force compliance with ranking transparency requirements. This, in turn, may lead to more balanced and fair markets for business and corporate website users, as well as more competition between platforms, as intended by the EU legislator. These benefits contribute to the United Nations Sustainable Development Goal 10.3 (of Economic and Affairs, 2023), aimed at reducing outcome inequalities and promoting equal opportunities.
@@
Finally, our research explores approaches that can measure the quality of explanations provided about software and algorithmic processes vis-à-vis a legal standard, which is actually a crucial field of research. This is particularly due to the rapid development of i) machine learning techniques and tools, ii) methods to explain machine learning systems and their outputs (Speith, 2022), and iii) legal requirements imposing explanations when such systems are used (Hacker and Passoth, 2020). In the context of our paper, it is clear that the ranking systems at stake include machine learning technology, which can be difficult to explain to laypersons. As lawmakers impose on online platforms to disclose information about these systems, it is particularly beneficial to society to be able to measure the quality of the explanations provided and determine when an explanation is insufficient or of poor quality. Besides, this reasoning and the tools that we develop can be transposed to other contexts implying the compulsory provision of explanations in relation to machine learning tools or outputs, such as the forthcoming AI Act (see, for instance, Art. 13 of the proposed AI Act (European Commission and Technology, 2021)).
9. Conclusions & Future Directions
The EU’s P2B Regulation requires online platforms to clearly explain their ranking procedures. In this context, we studied how platforms like Amazon, Booking, Google, Bing, Tripadvisor, and Yahoo comply with this regulation.
With three legal experts and feedback from 130 users, we analyzed the documentation. Our findings showed varying compliance levels. One reason might be the lack of standardized assessment criteria or automated tools.
To address this, we introduced two automated assessment tools. First, given its popularity, we considered ChatGPT as an assessment method. Next, we proposed a hybrid method (designed to better handle long documentation) merging ChatGPT with answer retrieval techniques and a new metric, DoX, to measure documentation explainability.
When we compared the results from our automated tools to manual assessments, they agreed with human evaluations 63.88% of the time. This suggests these tools could help make the review process more efficient. Importantly, the method based on DoX was particularly good at assessing lengthy documents. Furthermore, its explanatory scores aligned well with expert opinions, implying that content with weak explanations tends to receive lower compliance ratings from professionals.
While these tools have limitations, their continuous assessment capability can be invaluable for regulators and the public. Combining them with expert judgment can promote transparent documentation, essential for a fair digital business space. This transparency aligns with goals like the UN’s Sustainable Development Goal 10.3, fostering an informed and equitable online ecosystem.
Future improvements could refine these tools through better prompt engineering, tailoring models like ChatGPT for compliance tasks, and including multimedia analysis capabilities.
Acknowledgements.
@@
F. Sovrano and A. Bacchelli acknowledge the support of the Swiss National Science Foundation for the SNF Project 200021_197227. M. Lognoul aknowledges the support of the European Union H2020 research and innovation program, Grant Agreement No. 958339 (DENiM).
Author Contributions
@@
F. Sovrano and M. Lognoul: conceptualization, methodology, data curation, original draft preparation, investigation, validation, formal analysis. F. Sovrano: software, visualization. A. Bacchelli: methodology conceptualization, review, and editing.